A certain transverse wave is described by

y(x,t)=Bcos[2π(xLtτ)],

where B = 6.00 mm , L = 30.0 cm , and τ = 3.10×10−2 s .

Determine the wave's amplitude. A??

Determine the wave's wavelength.

λ??

Determine the wave's frequency.

f???

Determine the wave's speed of propagation.

v???
Answer #1

Here standard equation is:

= L = 30.0 cm

Part A.

Amplitude = factor by which cosine function is multiplied

Amplitude = B = 6.00 mm = 6.00*10^-3 m

Part B.

wavelength =

from above given values:

= 30.0 cm = 0.30 m = wavelength

Part C.

frequency = f = 1/

given that = 3.10*10^-2 sec

f = 1/(3.10*10^-2) = 32.258 Hz

f = 32.3 Hz = frequency

Part D.

wave's speed of propagation is

v = f*

v = 32.258*0.30 = 9.6774 m/sec

v = 9.68 m/sec

Please Upvote.